For what values of x will the graph of 2x - 3y + 5 = 0 be above the x-axis?
Answer by Alan3354(69443)

For what values of x will the graph of 2x - 3y + 5 = 0 be above the x-axis?
--------------
For all values of x which give a value of y >0. Set y = 0:
2x + 5 = 0
x = -5/2
---------
At x = -5/2, the line crosses the x-axis. For all values of x>(-5/2) it is above the x-axis.